# Redefining the Shortest Path Problem Formulation of the Linear Non-Gaussian Acyclic Model: Pairwise Likelihood Ratios, Prior Knowledge, and Path Enumeration

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: yyyydoi: XXXXXXX.XXXXXXXjournal: TKDDjournalvolume: vvjournalnumber: narticle: aaapublicationmonth: 01ccs: Computing methodologies Causal reasoning and diagnosticsccs: Mathematics of computing Causal networksccs: Information systems Data mining
@@ 1. Introduction
    Incorporating Known Relative Ordering: We expand LiNGAM-SPP to incorporate prior knowledge, requiring only relative ordering for better adaptability.
@@ 2.3. LiNGAM-SPP
3. Methods
In this section, we present the implementation details of our contributions. We begin by discussing the enhancement of LiNGAM-SPP by using PLR as its mutual information estimator. Subsequently, we show how a constrained path search, which involves skipping certain nodes, can be used to incorporate prior knowledge of relative orderings. Finally, we discuss our approach for predicting causal graph properties from the path distribution.
3.1. Improving LiNGAM-SPP with PLR
The initial LiNGAM-SPP paper (Suzuki and Yang, 2022) demonstrated that using the $k$NN-based mutual information estimator (Kraskov et al., 2004) with LiNGAM-SPP led to superior performance compared to the HSIC-based DirectLiNGAM. However, a limitation was that the parameter $k$ had to be tuned to achieve optimal performance. A subsequent study by the same authors (Suzuki and Yang, 2024) used copula entropy (Ma and Sun, 2011) as the mutual information estimator for LiNGAM-SPP, which, while achieving commendable performance, still relied on the $k$NN-based mutual information estimator (Kraskov et al., 2004). In this study, we propose using the PLR as an alternative to the $k$NN-based mutual information estimator. Notably, PLR eliminates the need for parameter tuning, demonstrates superior performance, and is computationally more efficient.
PLR uses the likelihood ratio, or the difference between the log-likelihoods, to determine the causal direction between two non-Gaussian variables. For instance, in determining the causal direction between $x$ and $y$, both standardized to have a mean of 0 and a variance of 1, PLR computes the difference between the log-likelihoods of both causal directions:
(7) $R=\frac{1}{T}\log L(x\rightarrow y)-\frac{1}{T}\log L(y\rightarrow x)$
Here, the log-likelihoods are normalized by the sample size $T$. Given $R$, we can infer the correct direction: if $R$ is positive, then $x\rightarrow y$; otherwise, $y\rightarrow x$. If $x\rightarrow y$, the SCM is $y=\rho x+d$, where $d\perp\!\!\!\perp x$. On the other hand, if $y\rightarrow x$, then the SCM is $x=\rho y+e$, where $e\perp\!\!\!\perp y$. Note that the coefficient $\rho$ is the same for both directions as it represents the correlation coefficient. Normally, computing R requires knowing the distributions of $x,y,d,e$. However, Hyvärinen (2013) (Hyvärinen and Smith, 2013) derived a practical approximation to compute this without prior knowledge of the distributions using differential entropy:
(8) $R\longrightarrow-H(x)-H(\hat{d}/\sigma_{d})+H(y)+H(\hat{e}/\sigma_{e})$
Here, $\hat{d}=y-\rho x$ and $\hat{e}=x-\rho y$ represent the estimated residuals, and $\sigma_{d}$ and $\sigma_{e}$ denote their standard deviations. $H$ represents the differential entropy, whose closed-form can be approximated using the approach by Hyvärinen (1998) (Hyvärinen, 1998):
(9) $\hat{H}(u)=H(v)-k_{1}[E\{\log\cosh u\}-\gamma]^{2}-k_{2}[E\{u\exp(-u^{2}/2)\}]^{2}$
where $H(v)=\frac{1}{2}(1+\log 2\pi)$, and the constants have numerical values $k_{1}\approx 79.047$, $k_{2}\approx 7.4129$, and $\gamma\approx 0.37457$. This closed-form approximation contributes significantly to the efficiency of PLR.
In our implementation, we incorporated code from the lingam package (Ikeuchi et al., 2023), which currently defaults to using PLR in its DirectLiNGAM implementation. We adapted the relevant portions of this code to integrate the PLR measure into the LiNGAM-SPP framework with slight modifications. Specifically, the original implementation calculated the PLR independence measure $m_{i}$ for each feature $i$ as follows:
(10) $m_{i}=-\sum_{j}\min(0,[M]_{ij})^{2}$
where $[M]_{ij}$ is the PLR between feature $i$ and $j$ (Hyvärinen and Smith, 2013). We introduce two modifications to adapt it to the shortest-path framework: First, we removed the negative sign and reformulated it as a minimization problem rather than a maximization one. Second, we normalized $m_{i}$ by dividing it by the number of terms in the summation. This adjustment is crucial because failing to do so could result in earlier steps having larger values of $m_{i}$ because they involve more terms. Such an imbalance in path measurements might pose challenges in identifying the shortest path.
3.2. Incorporating Known Relative Ordering
One notable feature of DirectLiNGAM is its capability to incorporate prior knowledge to enhance overall results by allowing users to specify which directed edges are required, forbidden, or unknown (Shimizu et al., 2011). In this study, we introduce an enhancement to LiNGAM-SPP to allow for integrating prior knowledge. Unlike DirectLiNGAM, where the exact edges need to be specified, our implementation only requires specifying the relative order. In practical terms, this means we can assert that one feature causes another while still allowing for the possibility of having intermediary features, or mediators, between them.
To implement this, we modify the shortest path algorithm to selectively skip nodes that violate the specified order. The path formulation of LiNGAM-SPP naturally lends itself to this task. Each node in the graph (e.g., Figure 1) contains information about relative ordering, indicating which features have been selected and which are yet to be explored. Specifically, if a feature is absent from the subscripts, it is positioned earlier in the causal order. For instance, referencing the same graph (Figure 1), if we assert that $x_{1}$ precedes $x_{2}$, then nodes with subscript 11{1}1 but not 22{2}2—such as $\{r_{3}^{(2)},r_{1}^{(2)}\}$ and $\{r_{1}^{(2,3)}\}$—should be skipped. This is because they imply that $x_{2}$ comes before $x_{1}$, which violates the known ordering.
3.3. Predicting Causal Graph Properties from Path Distributions
The utility of the path formulation extends beyond finding the shortest path. We discovered that path enumeration, representing the distribution of total path lengths of all possible orderings, reveals crucial insights into graph properties. Specifically, we observed and will demonstrate that the path distribution can be used to infer graph properties such as the presence of unmeasured confounders (an essential assumption to verify, especially for LiNGAM), the sparsity of the graph, and, to some extent, the reliability of PLR DirectLiNGAM and PLR LiNGAM-SPP.
We have observed, and will show in the results section, that clear patterns emerge in the path distributions in different cases, i.e. whether or not there are unmeasured confounders and differing graph sparsity. To take advantage of this insight – that the path distribution informs us about graph properties – we extract standardized moments to capture the shape of the distribution. Specifically, we compute the standardized moments of the log-transformed path lengths. These standardized moments will then serve as features in models trained to predict the various graph properties. The standardized moment is defined as follows:
(11) $\tilde{\mu}_{k}=\frac{\text{E}[(X-\mu)^{k}]}{(\text{Var}[X-\mu])^{k/2}}$
Here, $\tilde{\mu}_{k}$ denotes the $k$-th standardized moment, and $X$ represents the log-transformed path lengths.
After extracting the standardized moments, we train the following models using various popular machine learning algorithms ($k$NN (Pedregosa et al., 2011), Random Forest (Pedregosa et al., 2011), XGBoost (Chen and Guestrin, 2016), CatBoost (Dorogush et al., 2018), and AdaBoost (Pedregosa et al., 2011)):
@@ 3.3. Predicting Causal Graph Properties from Path Distributions
    Confounder Detector - a classifier to detect the presence of unmeasured confounders
@@
    Sparsity Estimator - a classifier and a regression model to estimate graph sparsity
@@
    DirectLiNGAM and LiNGAM-SPP Performance Predictors - classifier and regression models to predict the reliability of these methods
Although enumerating all possible paths can be done through brute force, we opted for a more efficient approach using Zero-Suppressed Binary Decision Diagrams (ZDDs). ZDDs are compact data structures that efficiently represent a family (or set) of sparse subsets from a finite ordered universe $U$ (Toda et al., 2017; Minato, 1993). In the context of graph theory, the diagram may also represent a family of subgraphs $G^{\prime}\subseteq G$ for a graph $G=(V,E)$. The frontier-based search algorithm provides a way to construct the diagram representing subgraphs, particularly paths, in a graph by setting the universe as the set of edges $U=E$ (Kawahara et al., 2017).
While the TdZdd library is commonly used for ZDD construction and set operations, its C++ implementation posed compatibility issues with our Python pipeline. Hence, we adopted the graphillion library instead, which offers a Python interface for ZDD operations. Although the standardized moments were calculated post hoc in our implementation, direct computation using ZDD operations, as demonstrated by (Lim et al., tted), is also possible.
4. Results
4.1. Improving LiNGAM-SPP with PLR
As previously discussed, we propose using PLRs as the mutual information estimator for LiNGAM-SPP, which we refer to as PLR LiNGAM-SPP. To evaluate the performance of PLR LiNGAM-SPP, we conducted experiments on a data generating process adapted from the DirectLiNGAM paper (Shimizu et al., 2011), with modifications to incorporate simulated unmeasured confounders. This data generating process differs from, but is more comprehensive than, the one utilized in the original LiNGAM-SPP studies (Suzuki and Yang, 2022, 2024), so the results here may be different from these studies. Furthermore, we assessed the method’s efficacy in practical settings by testing it on real-world datasets, comparing the results with those obtained from DirectLiNGAM. This validation demonstrates the method’s utility beyond simulated scenarios.
4.1.1. Data Generating Process
Our data generating process was adapted from (Shimizu et al., 2011), which we extended to incorporate cases with unmeasured confounders, following the framework outlined in (Tashiro et al., 2013; Hoyer et al., 2008), where LiNGAM with unmeasured confounders can be represented as:
(12) $\mathbf{x}=\mathbf{Bx}+\boldsymbol{\Lambda}\mathbf{f}+\mathbf{e}$
In this equation, $\mathbf{x}$ represents the observed data, $\mathbf{B}$ is a $p\times p$ matrix where $p$ is the number of features, $\mathbf{f}$ is the latent confounder vector, and $\mathbf{e}$ is the non-Gaussian noise. The matrix $\boldsymbol{\Lambda}$, with elements $\lambda_{ij}$, denotes the connection strength between $f_{j}$ and $x_{i}$. Additionally, each unmeasured confounder should have at least 2 children, and the matrix $\boldsymbol{\Lambda}$ should be of full column rank (Tashiro et al., 2013; Hoyer et al., 2008). To implement this, we performed the following steps:
@@ 4.1.1. Data Generating Process
    We initialized the matrix $\mathbf{B}$ by creating a $p\times p$ lower triangular matrix with elements randomly drawn from a uniform distribution in the interval $[-1.5,-0.5]\cup[0.5,1.5]$, following (Silva et al., 2006; Shimizu et al., 2011).
@@
    For modeling non-Gaussian noise and confounders, we used the set of 18 non-Gaussian distributions used in (Bach and Jordan, 2003; Shimizu et al., 2011). Each noise term and confounder were randomly sampled from this set. Moreover, we scaled these distributions to ensure their variances fell within the interval $[1,3]$, as in (Silva et al., 2006; Shimizu et al., 2011). In addition, we included a scaling parameter for the confounders called ”confounding strength”.
@@
    To construct $\boldsymbol{\Lambda}$, we generated its columns sequentially. For each column to be added, we first selected two variables to be influenced. Subsequently, we created a binary column vector by sampling it from a Bernoulli distribution, where the probability of success, an adjustable parameter, may be interpreted as the ”confoundedness” of the graph. The two selected variables and the binary vector are combined to create a vector determining which features are to be confounded. In effect, each column of $\boldsymbol{\Lambda}$ will have at least 2 nonzero elements, even if ”confoundedness” is set to 0. Finally, before appending the column to $\boldsymbol{\Lambda}$, we make sure that it is not equal to any of the existing columns, ensuring that $\boldsymbol{\Lambda}$ remains full rank.
@@
In summary, our data generating process is designed to capture a broad range of linear non-Gaussian processes. It allows us to control various parameters such as the number of features $p$, sample size $N$, sparsity, confoundedness, confounding strength, and the number of confounders. In the following sections where we present results of simulations, the simulations used sparsity drawn from a uniform distribution $\text{U}(0,1)$, confounding strength $10^{s}$ with $s$ drawn from $\text{U}(1,2)$, the number of confounders (if present) is randomly selected from $\{1,2,3\}$, and confoundedness is drawn from $\text{U}(0,1)$.
4.1.2. Comparing PLR LiNGAM-SPP with Existing Methods
@@
To showcase the effectiveness of this enhancement, we compare our proposed method, PLR LiNGAM-SPP, with other existing methods: PLR DirectLiNGAM (Hyvärinen and Smith, 2013), $k$NN LiNGAM-SPP (Suzuki and Yang, 2022), and copent (copula entropy) LiNGAM-SPP (Suzuki and Yang, 2024). Following the methodology in the original $k$NN LiNGAM-SPP paper (Suzuki and Yang, 2022), we used the following definition for the $k$NN-based mutual information estimator (Kraskov et al., 2004):
(13) $I(X,Y)=\psi(k)-\big{<}\psi(n_{x}+1)+\psi(n_{y}+1)\big{>}+\psi(N)$
@@
where $\psi(.)$ represents the digamma function. To implement this, we adapted code from the feature selection module of the sklearn (Pedregosa et al., 2011) library, extending it to accomodate a one-to-many variable setting, i.e. $I((X_{1},X_{2},\ldots,X_{m-1}),X_{m})$, by redefining $n_{x}$ in Eq. 13 to represent the number of points in the $(m-1)$-dimensional space, as outlined by (Kraskov et al., 2004). Meanwhile, we use the Python package copent (Ma and Sun, 2011) for computing the copula entropy in copent LiNGAM-SPP. Finally, performance was evaluated using the same error metric as previous LiNGAM-SPP papers (Suzuki and Yang, 2022, 2024), where the ordering error, denoted as $E_{o}$, is defined by:
@@
where $p$ represents the number of features, and $r$ is the count of pairs in the wrong order. This essentially counts the fraction pairs in the incorrect order.
Table 1. Comparison of Average Ordering Errors ($E_{o}$) Across Various Methods and Scenarios
@@
5 PWLR LiNGAM-SPP 0.29 0.28 0.23 0.21 0.46 0.45 0.42 0.39
PWLR DirectLiNGAM 0.29 0.28 0.23 0.22 0.46 0.45 0.42 0.40
@@
10 PWLR LiNGAM-SPP 0.33 0.24 0.21 0.18 0.44 0.40 0.37 0.34
PWLR DirectLiNGAM 0.32 0.25 0.22 0.19 0.44 0.41 0.37 0.36
@@
15 PWLR LiNGAM-SPP 0.33 0.24 0.19 0.16 0.41 0.36 0.35 0.29
PWLR DirectLiNGAM 0.32 0.25 0.21 0.18 0.42 0.37 0.33 0.32
@@
As previously mentioned, both $k$NN and copent LiNGAM-SPP require the tuning of the $k$ parameter, which is not possible in practical settings. However, for the purpose of comparison, instead of parameter tuning, we set $k$ to 5%, 10%, and the square root of the sample size ($N$), then selected the best-performing configuration. Table 1 presents results for various scenarios. The reported values represent the average ordering errors, $E_{o}$, across 250 trials for each scenario. The results and conclusions drawn from this comparison may differ from the previous LiNGAM-SPP studies, since a different, albeit more general, data generating process was used for the simulations.
While $k$NN LiNGAM-SPP excels for $p=5$, PLR LiNGAM-SPP outperforms in most cases for larger $p$, with DirectLiNGAM closely trailing. A paired t-test reveals that, overall, PLR LiNGAM-SPP shows a statistically significant improvement over DirectLiNGAM with a p-value of $5.22\times 10^{-5}$.
Interestingly, the performance disparity between PLR LiNGAM-SPP and DirectLiNGAM widens as the number of features ($p$) increases. We hypothesize that LiNGAM-SPP demonstrates greater robustness, particularly with longer sequences, owing to its ability for backtracking and correction. In contrast to DirectLiNGAM, which is prone to error propagation—wherein a mistake in one step compromises subsequent steps—LiNGAM-SPP, utilizing Dijkstra’s algorithm, exhibits a mechanism for error correction: If an incorrect feature is selected in one step, subsequent steps are more likely to yield large mutual information values. In such instances, Dijkstra’s algorithm can backtrack to explore alternative paths with smaller overall values.
Table 2. Average Runtime Comparison Across Various Methods and Scenarios (Note: m represents minutes)
@@
p MI Estimator
5 PWLR LiNGAM-SPP 0.44s 0.44s 0.45s 0.46s 0.43s 0.45s 0.44s 0.46s
PWLR DirectLiNGAM 0.04s 0.04s 0.04s 0.05s 0.04s 0.04s 0.04s 0.05s
@@
10 PWLR LiNGAM-SPP 6.8s 6.9s 6.9s 6.1s 6.4s 7.2s 7s 7.6s
PWLR DirectLiNGAM 0.26s 0.27s 0.3s 0.36s 0.26s 0.27s 0.31s 0.36s
@@
15 PWLR LiNGAM-SPP 2.2m 1.6m 1.2m 1.5m 1.4m 3.6m 2.8m 3.2m
PWLR DirectLiNGAM 0.83s 0.9s 0.98s 1.2s 0.84s 0.92s 1s 1.2s

### v2

@@ Abstract.
ccs: Computing methodologies Causal reasoning and diagnosticsccs: Mathematics of computing Causal networksccs: Information systems Data mining
@@ 1. Introduction
    Incorporating Known Relative Ordering: We expand LiNGAM-SPP to incorporate prior knowledge, requiring only relative orderings for better adaptability.
@@
An overview of the paper is as follows. Preliminary concepts relevant to the study – graphical causality, LiNGAM, and LiNGAM-SPP – are contained in the second section. The succeeding three discussion sections cover the threefold contribution of the study. The modified LiNGAM-SPP and evidence of its relative superiority is laid out in the third section. The fourth section details how prior knowledge may be considered in the framework through the input and processing of known relative orderings. The section on predicting the performance of causal discovery algorithms follows, hinging on path enumeration through the use of a compact data structure called the zero-suppressed binary decision diagram (ZDD). Each of the aforementioned discussion sections presents both the methods and results, specifying employed techniques, key statistics, and relevant observations. Such an arrangement is put forward for ease of readability. The work is closed in the final section with a recapitulation if the contributions.
@@ 2.3. LiNGAM-SPP
3. Improving LiNGAM-SPP
In the initial LiNGAM-SPP paper, it was demonstrated that LiNGAM-SPP outperforms DirectLiNGAM, specifically when comparing the $k$NN-based LiNGAM-SPP and the HSIC-based DirectLiNGAM (Suzuki and Yang, 2022). A succeeding paper (Suzuki and Yang, 2024) then showed that the LiNGAM-SPP based on copula entropy (copent) outperforms different implementations of DirectLiNGAM – PLR (Hyvärinen and Smith, 2013), kernel (Shimizu et al., 2011), HSIC (Hyvärinen and Smith, 2013), and ICA (Shimizu et al., 2006). While this initial success is promising, relying on the $k$NN-based mutual information raises concerns because it requires tuning the parameter $k$ for optimal performance. Unfortunately, this parameter tuning is not feasible in practical applications where the true data generating process is unknown.
To better test the performance of the method, we aim to test LiNGAM-SPP on a more general class of simulations and on real-world data. Given the above objectives, we introduce the following improvements on which the subsections are based.
@@ 3. Improving LiNGAM-SPP
    We adopt the data generating process used in the DirectLiNGAM paper (Shimizu et al., 2011) with a few additions for simulating confounders. This process is more general than the one used in the previous LiNGAM-SPP papers, providing a broader scope for testing and evaluation.
@@
    Instead of the $k$NN-based mutual information estimator (Kraskov et al., 2004) for LiNGAM-SPP, we propose adopting the PLR as the independence measure (Hyvärinen and Smith, 2013). This modification not only enhances performance and computation time but also eliminates the need for fine-tuning.
@@
    To validate the effectiveness of the method in practical settings, we conduct tests on LiNGAM-SPP using real-world data and compared its results with DirectLiNGAM. This step provides insights into the applicability of the method beyond simulated scenarios.
3.1. Data Generating Process
Our data generating process was adapted from (Shimizu et al., 2011), which we extended to incorporate cases with unmeasured confounders. The framework outlined in (Tashiro et al., 2013; Hoyer et al., 2008) is followed, where LiNGAM with unmeasured confounders can be represented as
(7) $\mathbf{x}=\mathbf{Bx}+\boldsymbol{\Lambda}\mathbf{f}+\mathbf{e}.$
In the equation above, $\mathbf{x}$ represents the observed data, $\mathbf{B}$ is a $p\times p$ matrix where $p$ is the number of features, $\mathbf{f}$ is the latent confounder vector, and $\mathbf{e}$ is the non-Gaussian noise. The matrix $\boldsymbol{\Lambda}$, with elements $\lambda_{ij}$, denotes the connection strength between $f_{j}$ and $x_{i}$. Additionally, each unmeasured confounder should have at least 2 children and the matrix $\boldsymbol{\Lambda}$ should be of full column rank (Tashiro et al., 2013; Hoyer et al., 2008). To implement this, we perform the following steps.
@@ 3.1. Data Generating Process
    We initialize the matrix $\mathbf{B}$ by creating a $p\times p$ lower triangular matrix with elements randomly drawn from a uniform distribution in the interval $[-1.5,-0.5]\cup[0.5,1.5]$, following (Silva et al., 2006; Shimizu et al., 2011).
@@
    For modeling non-Gaussian noise and confounders, we used the set of 18 non-Gaussian distributions used in (Bach and Jordan, 2003; Shimizu et al., 2011). Each noise term and confounder were randomly sampled from this set. Moreover, we scaled these distributions to ensure that their variances fell within the interval $[1,3]$, as in (Silva et al., 2006; Shimizu et al., 2011). We also included a scaling parameter for the confounders which we refer to as confounding strength.
@@
    To construct $\boldsymbol{\Lambda}$, we generated its columns sequentially. For each column to be added, we first selected two variables to be influenced. Subsequently, we created a binary column vector by sampling it from a Bernoulli distribution where the probability of success, an adjustable parameter, may be interpreted as the confoundedness of the graph. The two selected variables and the binary vector are combined to create a vector determining which features are to be confounded. In effect, each column of $\boldsymbol{\Lambda}$ will have at least 2 nonzero elements, even if the confoundedness is set to 0. Before appending the column to $\boldsymbol{\Lambda}$, we make sure that it is not equal to any of the existing columns, hence guaranteeing that $\boldsymbol{\Lambda}$ remains full rank.
@@
Our data generating process is designed to capture a broad range of linear non-Gaussian processes. It allows us to control various parameters such as the number of features $p$, sample size $N$, sparsity, confoundedness, confounding strength, and the number of confounders. In the following sections where we present results of simulations, the simulations used sparsity drawn from a uniform distribution $\text{U}(0,1)$, confounding strength $10^{s}$ with $s$ drawn from $\text{U}(1,2)$, the number of confounders randomly drawn from $\{1,2,3\}$, and confoundedness drawn from $\text{U}(0,1)$.
3.2. PLR as an Independence Measure
The initial LiNGAM-SPP paper (Suzuki and Yang, 2022) demonstrated that using the $k$NN-based mutual information estimator (Kraskov et al., 2004) with LiNGAM-SPP led to superior performance compared to the HSIC-based DirectLiNGAM. However, a limitation was that the parameter $k$ had to be tuned to achieve optimal performance. A subsequent study by the same authors (Suzuki and Yang, 2024) used copent (Ma and Sun, 2011) as the mutual information estimator for LiNGAM-SPP which, while achieving commendable performance, still relied on the $k$NN-based mutual information estimator (Kraskov et al., 2004). In this study, we propose using the PLR as an alternative to the $k$NN-based mutual information estimator. Notably, PLR eliminates the need for parameter tuning, demonstrates superior performance, and is computationally more efficient.
PLR uses the likelihood ratio, the difference between the log-likelihoods, to determine the causal direction between two non-Gaussian variables. For instance, in determining the causal direction between $x$ and $y$, both standardized to have a mean of 0 and a variance of 1, PLR computes the difference between the log-likelihoods of both causal directions
(8) $R=\frac{1}{T}\log L(x\rightarrow y)-\frac{1}{T}\log L(y\rightarrow x).$
Here, the log-likelihoods are normalized by the sample size $T$. Given $R$, we can infer the correct direction. If $R$ is positive, then $x\rightarrow y$; otherwise, $y\rightarrow x$. If $x\rightarrow y$, the SCM is $y=\rho x+d$, where $d\perp\!\!\!\perp x$. On the other hand, if $y\rightarrow x$, then the SCM is $x=\rho y+e$, where $e\perp\!\!\!\perp y$. Note that the coefficient $\rho$ is the same for both directions as it represents the correlation coefficient. Normally, computing $R$ requires knowing the distributions of $x$, $y$, $d$, and $e$. A practical approximation using differential entropy (Hyvärinen and Smith, 2013) was nonetheless derived to compute this without prior knowledge of the distributions. Mathematically,
(9) $R\longrightarrow-H(x)-H(\hat{d}/\sigma_{d})+H(y)+H(\hat{e}/\sigma_{e}).$
Here, $\hat{d}=y-\rho x$ and $\hat{e}=x-\rho y$ represent the estimated residuals and $\sigma_{d}$ and $\sigma_{e}$ denote their standard deviations. The closed-form of the differential entropy $H$ may be approximated by
(10) $\hat{H}(u)=H(v)-k_{1}[E\{\log\cosh u\}-\gamma]^{2}-k_{2}[E\{u\exp(-u^{2}/2)\}]^{2},$
where $H(v)=\frac{1}{2}(1+\log 2\pi)$ and the constants have numerical values $k_{1}\approx 79.047$, $k_{2}\approx 7.4129$, and $\gamma\approx 0.37457$ (Hyvärinen, 1998). This closed-form approximation contributes significantly to the efficiency of PLR.
In our implementation, we incorporate code from the lingam package (Ikeuchi et al., 2023), which currently defaults to using PLR in its DirectLiNGAM implementation. We use the relevant portions of this code to integrate the PLR measure into the LiNGAM-SPP framework with slight modifications. Specifically, the original implementation calculated the PLR independence measure $m_{i}$ for each feature $i$ as
(11) $m_{i}=-\sum_{j}\min(0,[M]_{ij})^{2},$
where $[M]_{ij}$ is the PLR between feature $i$ and $j$ (Hyvärinen and Smith, 2013). We introduce two modifications to adapt it to the shortest-path framework. First, we remove the negative sign and reformulated it as a minimization problem rather than a maximization one. Second, we normalize $m_{i}$ by dividing it by the number of terms in the summation. This adjustment is crucial because failing to do so could result in earlier steps having larger values of $m_{i}$ because they involve more terms. Such an imbalance in path measurements might pose challenges in identifying the shortest path.
As previously discussed, we propose using the PLR as the mutual information estimator for LiNGAM-SPP, which we refer to as PLR LiNGAM-SPP. To evaluate the performance of PLR LiNGAM-SPP, we conducted experiments on a data generating process adapted from the DirectLiNGAM paper (Shimizu et al., 2011), with modifications to incorporate simulated unmeasured confounders. This data generating process differs from, but is more comprehensive than, the one utilized in the original LiNGAM-SPP studies (Suzuki and Yang, 2022, 2024), so the results here may be different from these studies. Furthermore, we assessed the efficacy of the method in practical settings by testing it on real-world datasets, comparing the results with those obtained from DirectLiNGAM. This validation demonstrates the utility of the method beyond simulated scenarios.
@@ 3.2. PLR as an Independence Measure
To showcase the effectiveness of this enhancement, we compare our proposed method, PLR LiNGAM-SPP, with other existing methods – PLR DirectLiNGAM (Hyvärinen and Smith, 2013), $k$NN LiNGAM-SPP (Suzuki and Yang, 2022), and copent LiNGAM-SPP (Suzuki and Yang, 2024). Following the methodology in the original $k$NN LiNGAM-SPP paper (Suzuki and Yang, 2022), we used(Kraskov et al., 2004)
(12) $I(X,Y)=\psi(k)-\big{<}\psi(n_{x}+1)+\psi(n_{y}+1)\big{>}+\psi(N),$
@@
where $\psi(.)$ represents the digamma function, as the definition for the $k$NN-based mutual information estimator. To implement this, we adapted code from the feature selection module of the sklearn (Pedregosa et al., 2011) library. We extend it to accomodate a one-to-many variable setting, i.e., $I((X_{1},X_{2},\ldots,X_{m-1}),X_{m})$, by redefining $n_{x}$ in Eq. 12 to represent the number of points in the $(m-1)$-dimensional space as outlined by (Kraskov et al., 2004). Meanwhile, we use the Python package copent (Ma and Sun, 2011) for computing the copent in copent LiNGAM-SPP. Finally, performance was evaluated using the same error metric as previous LiNGAM-SPP papers (Suzuki and Yang, 2022, 2024) where the ordering error, denoted as $E_{o}$, is defined by
@@
where $p$ represents the number of features and $r$ is the count of pairs in the wrong order. This essentially counts the fraction pairs in the incorrect order.
As previously mentioned, both $k$NN and copent LiNGAM-SPP require the tuning of the $k$ parameter, which is not possible in practical settings. However, for the purpose of comparison, we set $k$ to 5%, 10%, and the square root of the sample size ($N$) instead of parameter tuning. We then select the best-performing configuration. Table 1 presents results for various scenarios. The reported values represent the average ordering errors, $E_{o}$, across 250 trials for each scenario. The results and conclusions drawn from this comparison may differ from the previous LiNGAM-SPP studies since a different, albeit more general, data generating process was used for the simulations.
Table 1. Average Ordering Errors $E_{o}$ Across Various Methods and Scenarios
@@
5 PLR LiNGAM-SPP 0.29 0.28 0.23 0.21 0.46 0.45 0.42 0.39
PLR DirectLiNGAM 0.29 0.28 0.23 0.22 0.46 0.45 0.42 0.40
@@
10 PLR LiNGAM-SPP 0.33 0.24 0.21 0.18 0.44 0.40 0.37 0.34
PLR DirectLiNGAM 0.32 0.25 0.22 0.19 0.44 0.41 0.37 0.36
@@
15 PLR LiNGAM-SPP 0.33 0.24 0.19 0.16 0.41 0.36 0.35 0.29
PLR DirectLiNGAM 0.32 0.25 0.21 0.18 0.42 0.37 0.33 0.32
@@
While $k$NN LiNGAM-SPP excels for $p=5$, PLR LiNGAM-SPP outperforms in most cases for larger $p$ with DirectLiNGAM closely trailing. A paired t-test reveals that overall, PLR LiNGAM-SPP shows a statistically significant improvement over DirectLiNGAM with a p-value of $5.22\times 10^{-5}$.
Interestingly, the performance disparity between PLR LiNGAM-SPP and DirectLiNGAM widens as the number of features $p$ increases. We hypothesize that LiNGAM-SPP demonstrates greater robustness particularly with longer sequences, owing to its ability for backtracking and correction. In contrast to DirectLiNGAM which is prone to error propagation – wherein a mistake in one step compromises subsequent steps, LiNGAM-SPP utilizing Dijkstra’s algorithm exhibits a mechanism for error correction. If an incorrect feature is selected in one step, subsequent steps are more likely to yield large mutual information values. In such instances, Dijkstra’s algorithm can backtrack to explore alternative paths with smaller overall values.
Table 2. Average Runtime Across Various Methods and Scenarios
@@
p Method
5 PLR LiNGAM-SPP 0.44s 0.44s 0.45s 0.46s 0.43s 0.45s 0.44s 0.46s
PLR DirectLiNGAM 0.04s 0.04s 0.04s 0.05s 0.04s 0.04s 0.04s 0.05s
@@
10 PLR LiNGAM-SPP 6.8s 6.9s 6.9s 6.1s 6.4s 7.2s 7s 7.6s
PLR DirectLiNGAM 0.26s 0.27s 0.3s 0.36s 0.26s 0.27s 0.31s 0.36s
@@
15 PLR LiNGAM-SPP 2.2m 1.6m 1.2m 1.5m 1.4m 3.6m 2.8m 3.2m
PLR DirectLiNGAM 0.83s 0.9s 0.98s 1.2s 0.84s 0.92s 1s 1.2s